# SceneSplat++: A Large Dataset and Comprehensive Benchmark for Language Gaussian Splatting

### v1

@@ Abstract
3D Gaussian Splatting (3DGS) serves as a highly performant and efficient encoding of scene geometry, appearance, and semantics. Moreover, grounding language in 3D scenes has proven to be an effective strategy for 3D scene understanding. Current Language Gaussian Splatting line of work falls into three main groups: (i) per-scene optimization-based, (ii) per-scene optimization-free, and (iii) generalizable approach. However, most of them are evaluated only on rendered 2D views of a handful of scenes and viewpoints close to the training views, limiting their ability and insight into holistic 3D understanding. To address this gap, we propose the first large-scale benchmark that systematically assesses these three groups of methods directly in 3D space, evaluating on 1060 scenes across three indoor datasets and one outdoor dataset. SceneSplat-Bench results demonstrate a clear advantage of the generalizable paradigm, particularly in relaxing the scene-specific limitation, enabling fast feed-forward inference on novel scenes, and achieving superior segmentation performance. We further introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising around 49K diverse indoor and outdoor scenes obtained from multiple sources, with which we demonstrate the generalizable approach could harness strong data priors. Our codes, benchmark, and datasets will be made public to accelerate research in generalizable 3DGS scene understanding.
@@ 1 Introduction
The field of 3D computer vision is highly focused on capturing the geometry WITKIN198117 ; Kazhdan2006 ; Hirschmuller2008StereoDepth ; Newcombe2011KinectFusion ; schoenberger2016sfm ; Park2019DeepSDF ; shi4568335city ; ren2024bringing and the visual appearance 10.5555/539405 ; 791235 ; 5226635 ; mildenhall2020nerf ; kerbl3Dgaussians of a scene, as well as understanding its content 8099499 ; wu2024ptv3 ; peng2023openscene ; zhu2023pointclip ; sarkar2025crossover . Recently, 3D Gaussian Splatting (3DGS) kerbl3Dgaussians has emerged as the most desirable 3D representation due to its unique ability to jointly encode geometry, appearance, and scene understanding properties Ye2024GSgroupping ; Shi2024LEG ; jiang2024open in a compact form that can be effectively optimized from 2D posed images ye2025gsplat . Furthermore, vision-language reasoning represents the most promising direction for 3D scene understanding rozenberszki2022language ; ding2022pla ; yang2024regionplc ; ma2025cityloc ; jiang2024open ; wu2023mars , as it bridges the visual and geometric attributes of a scene with the language we use to define, describe, and reason about concepts. Therefore, this work focuses exclusively on vision-language scene understanding with 3DGS.
@@
While these methods represent important advances in integrating vision-language reasoning with 3DGS, their evaluation protocols still suffer from several critical limitations. First, most methods are evaluated on only a small number of selected scenes Qin2024LangSplat ; Zuo2024FMGS ; zhou2024feature3DGS ; Wu2024OpenGaussians ; QU2024GOI ; joseph2024GWfeatureBP ; cheng2024occam . As a result, such evaluations carry a high risk of producing scene-specific rather than generalizable conclusions. Moreover, the absence of a standardized benchmark further limits systematic and comparable evaluation. Second, most methods are evaluated close to the training views Qin2024LangSplat ; marrie2024ludvig ; Wu2024OpenGaussians . Consequently, there is no guarantee that the results generalize well to novel viewpoints. Finally, most evaluations are conducted in 2D rather than rather than directly in 3D space Qin2024LangSplat ; Zuo2024FMGS ; zhou2024feature3DGS ; Wu2024OpenGaussians ; QU2024GOI ; joseph2024GWfeatureBP ; marrie2024ludvig ; cheng2024occam . However, 3D scene understanding fundamentally concerns performance in 3D space, which cannot be fully inferred from 2D projections. Therefore, we introduce SceneSplat-Bench, an evaluation benchmark for 3DGS vision-language scene understanding that includes 1,060 scenes spanning 325 unique semantic classes. It evaluates performance in 3D for each segment of the scene, and we use it to assess representative methods from all three groups, as presented in Tab. 1.
@@
Methods Generalizable Optimization- free Runtime / Scene Benchmark Dataset Scene Number Class Number
@@
Gradient-Weighted 3DGS  joseph2024GWfeatureBP 4 min LERF kerr2023lerf , 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
@@
SceneSplat-Bench ScanNet dai2017scannet , ScanNet++ yeshwanth2023scannet++ , Matterport3D chang2017matterport3d , HoliCity zhou2021holicity 1060 325
@@ 3 Dataset
Property SceneSplat-7K DL3DV-10K HoliCity Aria-ASE Crowdsourced Total
@@
Figure 2: Appearance, Geometry, and Scale Statistics of the SceneSplat-49K Dataset. Distributions of photometric (PSNR, SSIM, LPIPS) and geometric (depth $\ell_{1}$) reconstruction errors show consistently high-quality renders across scenes, while the wide spread in total Gaussian number and indoor/outdoor scene floor area demonstrates the dataset’s diversity. The curves are convolved from the bucket values and vertical dotted lines mark the mean of each metric.
@@ 3.2 Vision-Language Embedding Collection
Vision-language embeddings are extracted from the frames used to optimize the 3DGS. Unlike existing pretraining methods that align 3D primitives with text embeddings from vision-language models, we align each Gaussian directly in the image-embedding space, thereby preserving richer latent semantics. This approach avoids the information loss inherent in textual descriptions.
@@ 4 Benchmark
We formulate our benchmark as follows. A scene is represented by a set of $N$ Gaussian primitives $\mathcal{G}=\{G_{j}\}_{j=1}^{N}$, where each $G_{j}\in\mathbb{R}^{59}$ encodes its position, covariance, color, and opacity. For every primitive we compute a language embedding $E_{j}^{G}\in\mathbb{R}^{d}$, obtained by per-scene processing or by inference with a learned encoder. The user provides a set of free-form text queries $C=\{c_{i}\}_{i=1}^{n}$. A frozen text encoder $P_{T}$ (e.g., CLIP) maps each query to an embedding $E_{i}^{T}=P_{T}(c_{i})$. Each Gaussian is labeled with the query whose embedding is most similar to its own:
@@ 4.1 Benchmark settings
Evaluation Protocols. We evaluate 9 methods from 3 categories on our SceneSplat-Bench benchmark. We report two key metrics in 3D: foreground mean Intersection over Union (f-mIoU) and foreground mean accuracy (f-mAcc), which address object size imbalances and reduce viewpoint dependency compared with 2D evaluation. We unified the evaluation pipeline by first converting the language fields into a unified format $\mathcal{E}_{G}=\{E_{j}\}_{j=1}^{N}$, then computing the similarity with the queried text embedding and selecting the pair with the highest similarity. K-nearest neighbors are used to vote predictions from Gaussian splats to ground truth label positions. We exclude floor, wall, and ceiling classes as background for indoor scenes and the sky class for outdoor scenes. All runtime measurements were obtained on NVIDIA RTX A6000 GPUs, except for FMGS, whose training requires 50–80 GB of GPU memory; its runtime was therefore measured on NVIDIA H200 GPUs.
Benchmark Dataset. We evaluate the methods’ performance on indoor scenes using ScanNet (with 20 and 200 classes), ScanNet++ (100 classes), and Matterport3D (21 classes). Queries cover a wide range of object granularity—from small items like a mouse to larger objects such as a bed—reflecting the method’s capability to varying scales understanding. HoliCity (4 classes) is employed to evaluate the performance on outdoor scenes, emphasizing large structures under complex surrounding environment. Performance results are reported on randomly sampled 10-scene subsets for each of the four benchmarks across all methods. For methods demonstrating strong efficiency, we conduct a full evaluation on the complete validation set.
Baseline Methods. We categorize the methods based on their optimization paradigm and generalizability to novel scenes. (i) Per-Scene Optimization-based Methods assign learnable attributes to Gaussian splats and optimize them via backpropagation. However, rendering additional features with alpha compositing is computationally expensive and requires per-scene optimization. (ii) Per-Scene Optimization-Free methods use a training-free approach, lifting 2D features into 3D space as a weighted sum of 2D representations, requiring only a single forward rendering pass. (iii) Generalizable Method SceneSplat trains a feed-forward 3DGS encoder, learning to predict representations aligned with image embeddings. We further extend its training based on our dataset. SceneSplat (Pseudo Label) denotes the VLM embeddings used for its vision-language pretraining. We refer to the supplement for details on running each baseline.
@@
LangSplat 0.0153 0.0741 - - 0.0057 0.0106 - -
OpenGaussian 0.1186 0.2340 - - 0.0545 0.1086 - -
Feature-3DGS 0.1711 0.2512 - - 0.0541 0.0876 - -
FMGS 0.0916 0.2028 0.1169 0.2684 0.0641 0.1374 0.0579 0.1443
GOI 0.2295 0.4638 0.2115 0.4286 0.1193 0.2297 0.1001 0.2240
@@
Gradient-Weighted 3DGS 0.4184 0.5174 0.3573 0.4638 0.1638 0.2502 0.0601 0.1100
LUDVIG 0.1406 0.2885 0.1337 0.2969 0.0887 0.1833 0.0602 0.1701
OccamLGS 0.3559 0.4587 0.2308 0.4148 0.1936 0.2458 0.1204 0.2503
SceneSplat (Pseudo Label) 0.3521 0.5210 0.3500 0.5550 0.2520 0.3038 0.2280 0.3590
@@
SceneSplat 0.2968 0.5711 0.3540 0.5780 0.1387 0.4198 0.1648 0.3573
Table 3: Zero-shot 3D Semantic Segmentation Experiments on ScanNet20 (20 classes) and ScanNet200 (200 classes). All methods are evaluated on a 10-scene mini-validation set, with the 312-scene full validation set run only for selected methods due to runtime limitations (Tab. 1).
@@
LangSplat 0.0131 0.0332 - - 0.0219 0.1257 - -
OpenGaussian 0.0695 0.1186 - - 0.1387 0.2698 - -
Feature-3DGS 0.0798 0.1765 - - 0.3043 0.4732 - -
FMGS 0.0984 0.1892 0.1004 0.2300 0.2105 0.3664 - -
GOI 0.1593 0.2790 0.1631 0.3296 0.1631 0.3296 0.1652 0.3130
@@
Gradient-Weighted 3DGS 0.0925 0.1391 0.0900 0.1328 0.3083 0.4586 0.2762 0.4187
LUDVIG 0.0995 0.2099 0.1099 0.2547 0.2348 0.3888 0.1960 0.3738
OccamLGS 0.1580 0.2861 0.1502 0.3312 0.1803 0.3263 0.1725 0.3151
SceneSplat (Pseudo Label) 0.2340 0.3709 0.2243 0.4667 0.2392 0.4368 0.2748 0.4384
@@
SceneSplat 0.2263 0.4916 0.2836 0.4992 0.2732 0.5379 0.3384 0.5745
Table 4: Zero-Shot 3D Semantic Segmentation Experiments On ScanNet++ and Matterport3D. All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated only for selected methods due to runtime limitations (Tab. 1).
@@
Table 5: Impact of Training-Data Scaling on Indoor Benchmarks and Cross-Domain Generalization to HoliCity. Results show that more training data consistently improves indoor performance. Furthermore, models trained on indoor data only surprisingly transfer to outdoor scenes.
     Method      10 scenes      val set, 328 scenes
     f–mIoU      f–mAcc      f–mIoU      f–mAcc
     Per-Scene Optimization Methods
     LangSplat      0.0918      0.1950      -      -
     OpenGaussian      0.1853      0.2567      -      -
     Feature-3DGS      0.2479      0.5646      -      -
     FMGS      0.1774      0.2513      -      -
     GOI      0.1126      0.3773      0.1587      0.3772
     Per-Scene Optimization-Free Methods
     Gradient-Weighted 3DGS      0.2484      0.4808      0.2762      0.4187
     LUDVIG      0.1593      0.2296      0.1843      0.2775
     OccamLGS      0.2255      0.4754      0.2234      0.5056
     SceneSplat (Pseudo Label)      0.2464      0.4780      0.2522      0.5159
     Generalizable Method
     SceneSplat      0.3081      0.5647      0.2880      0.6045
Table 6: Zero-Shot 3D Outdoor Segmentation On HoliCity Dataset. All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated for selected methods.

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) serves as a highly performant and efficient encoding of scene geometry, appearance, and semantics. Moreover, grounding language in 3D scenes has proven to be an effective strategy for 3D scene understanding. Current Language Gaussian Splatting line of work falls into three main groups: (i) per-scene optimization-based, (ii) per-scene optimization-free, and (iii) generalizable approach. However, most of them are evaluated only on rendered 2D views of a handful of scenes and viewpoints close to the training views, limiting their ability and insight into holistic 3D understanding. To address this gap, we propose the first large-scale benchmark that systematically assesses these three groups of methods directly in 3D space, evaluating on 1060 scenes across three indoor datasets and one outdoor dataset. SceneSplat-Bench results demonstrate a clear advantage of the generalizable paradigm, particularly in relaxing the scene-specific limitation, enabling fast feed-forward inference on novel scenes, and achieving superior segmentation performance. We further introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising around 49K diverse indoor and outdoor scenes obtained from multiple sources, with which we demonstrate the generalizable approach could harness strong data priors. Our codes, benchmark, and datasets are available at SceneSplat-Bench and SceneSplat-49K.
@@ 1 Introduction
The field of 3D computer vision is highly focused on capturing the geometry WITKIN198117; Kazhdan2006; Hirschmuller2008StereoDepth; Newcombe2011KinectFusion; schoenberger2016sfm; Park2019DeepSDF; shi4568335city; ren2024bringing and the visual appearance 10.5555/539405; 791235; 5226635; mildenhall2020nerf; kerbl3Dgaussians of a scene, as well as understanding its content 8099499; wu2024ptv3; peng2023openscene; zhu2023pointclip; sarkar2025crossover. Recently, 3D Gaussian Splatting (3DGS) kerbl3Dgaussians has emerged as the most desirable 3D representation due to its unique ability to jointly encode geometry, appearance, and scene understanding properties Ye2024GSgroupping; Shi2024LEG; jiang2024open in a compact form that can be effectively optimized from 2D posed images ye2025gsplat. Furthermore, vision-language reasoning represents the most promising direction for 3D scene understanding rozenberszki2022language; ding2022pla; yang2024regionplc; ma2025cityloc; jiang2024open; wu2023mars; halacheva2025gaussianvlm, as it bridges the visual and geometric attributes of a scene with the language we use to define, describe, and reason about concepts. Therefore, this work focuses exclusively on vision-language scene understanding with 3DGS.
@@
While these methods represent important advances in integrating vision-language reasoning with 3DGS, their evaluation protocols still suffer from several critical limitations. First, most methods are evaluated on only a small number of selected scenes Qin2024LangSplat; Zuo2024FMGS; zhou2024feature3DGS; Wu2024OpenGaussians; QU2024GOI; joseph2024GWfeatureBP; cheng2024occam. As a result, such evaluations carry a high risk of producing scene-specific rather than generalizable conclusions. Moreover, the absence of a standardized benchmark further limits systematic and comparable evaluation. Second, most methods are evaluated close to the training views Qin2024LangSplat; marrie2024ludvig; Wu2024OpenGaussians. Consequently, there is no guarantee that the results generalize well to novel viewpoints. Finally, most evaluations are conducted in 2D rather than directly in 3D space Qin2024LangSplat; Zuo2024FMGS; zhou2024feature3DGS; Wu2024OpenGaussians; QU2024GOI; joseph2024GWfeatureBP; marrie2024ludvig; cheng2024occam. However, 3D scene understanding fundamentally concerns performance in 3D space, which cannot be fully inferred from 2D projections. Therefore, we introduce SceneSplat-Bench, an evaluation benchmark for 3DGS vision-language scene understanding that includes 1060 scenes spanning 325 unique semantic classes. It evaluates performance in 3D for each segment of the scene, and we use it to assess representative methods from all three groups, as presented in Tab.˜1.
@@
Methods
Generalizable
Optimization-
free
Runtime
/ Scene
Benchmark
Dataset
Scene
Number
Class
Number
@@ Methods
Gradient-Weighted
3DGS
 joseph2024GWfeatureBP
4 min LERF kerr2023lerf, 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
@@
SceneSplat-Bench
ScanNet dai2017scannet, ScanNet++ yeshwanth2023scannet++,
Matterport3D chang2017matterport3d, HoliCity zhou2021holicity,
1060 325
@@ 3 Dataset
Property SceneSplat‑7K li2025SceneSplat7k DL3DV‑10K ling2024dl3dv HoliCity zhou2021holicity Aria‑ASE avetisyan2024scenescript Crowdsourced Total
@@
Figure 2: Appearance, Geometry, and Scale Statistics of the SceneSplat-49K Dataset. Distributions of photometric (PSNR, SSIM, LPIPS) and geometric (depth $\ell_{1}$) reconstruction errors show consistently high‐quality renders across scenes, while the wide spread in total Gaussian number and indoor/outdoor scene floor area demonstrates the dataset’s diversity. The curves are convolved from the bucket values and vertical dotted lines mark the mean of each metric.
@@ 3.2 Vision-Language Embedding Collection
Vision-language embeddings are extracted from the frames used to optimize the 3DGS. Unlike existing pretraining methods that align 3D primitives with text embeddings from vision-language models, we align each Gaussian directly in the image‐embedding space, thereby preserving richer latent semantics. This approach avoids the information loss inherent in textual descriptions.
@@ 4 Benchmark
We formulate the assigning of Gaussian language feature as follows. A scene is represented by a set of $N$ Gaussian primitives $\mathcal{G}=\{G_{j}\}_{j=1}^{N}$, where each $G_{j}\in\mathbb{R}^{59}$ encodes its position, covariance, color, and opacity. For every primitive we compute a language embedding $E_{j}^{G}\in\mathbb{R}^{d}$, obtained by per-scene processing or by inference with a learned encoder. The user provides a set of free-form text queries $C=\{c_{i}\}_{i=1}^{n}$. A frozen text encoder $P_{T}$ (e.g., CLIP) maps each query to an embedding $E_{i}^{T}=P_{T}(c_{i})$. Each Gaussian is labeled with the query whose embedding is most similar to its own:
@@ 4.1 Benchmark settings
Evaluation Protocols. We evaluate 9 methods from 3 categories on our SceneSplat-Bench benchmark. We report two key metrics in 3D: foreground mean Intersection over Union (f-mIoU) and foreground mean accuracy (f-mAcc), which address object size imbalances and reduce viewpoint dependency compared with traditional 2D-only evaluation. We unified the evaluation pipeline by first converting the language fields into a unified format $\mathcal{E}_{G}=\{E_{j}\}_{j=1}^{N}$, then computing the cosine similarity between each language feature and the queried text embeddings. Each Gaussian primitive is assigned the semantic label corresponding to the most similar text embedding. For every ground-truth 3D point used for evaluation, we identify its K-nearest Gaussian neighbors (K=25) and assign the majority-voted label as the semantic prediction for that point. We exclude floor, wall, and ceiling classes as background for indoor scenes and the sky class for outdoor scenes. To make our evaluation more complete, we additionally evaluate semantics on 2D renders, following the protocol of LERF kerr2023lerf. The 2D evaluation is conducted on subsets of the benchmark datasets with two key 2D metrics: mean Intersection over Union (mIoU) and mean accuracy (mAcc) on all classes (background and foreground classes). All runtime measurements were obtained on NVIDIA RTX A6000 GPUs, except for FMGS, whose training requires 50–80 GB of GPU memory; its runtime was therefore measured on NVIDIA H200 GPUs.
Benchmark Dataset. We evaluate the methods’ performance on indoor scenes using ScanNet (with 20 and 200 classes), ScanNet++ (100 classes), and Matterport3D (21 classes). Queries cover a wide range of object granularity—from small items like a mouse to larger objects such as a bed—reflecting the method’s capability to varying scales understanding. HoliCity (4 classes) is employed to evaluate the performance on outdoor scenes, emphasizing large structures under complex surrounding environment. Performance results are reported on randomly sampled 10-scenes subsets for each of the four benchmarks across all methods. For methods demonstrating strong efficiency, we conduct a full evaluation on the complete validation set.
Baseline Methods. We categorize the methods based on their optimization paradigm and generalizability to novel scenes.
(i) Per-Scene Optimization-based Methods assign learnable attributes to Gaussian splats and optimize them via backpropagation. However, rendering additional features with alpha compositing is computationally expensive and requires per-scene optimization.
(ii) Per-Scene Optimization-Free methods use a training-free approach, lifting 2D features into 3D space as a weighted sum of 2D representations, requiring only a single forward rendering pass.
(iii) Generalizable Method SceneSplat trains a feed-forward 3DGS encoder, learning to predict representations aligned with image embeddings. We further extend its training based on our dataset. SceneSplat (Pseudo Label) denotes the VLM embeddings used for its vision-language pretraining. We refer to the supplement for details on running each baseline.
@@
LangSplat Qin2024LangSplat 0.0153 0.0741 - - 0.0057 0.0106 - -
OpenGaussian Wu2024OpenGaussians 0.1186 0.2340 - - 0.0545 0.1086 - -
Feature-3DGS zhou2024feature3DGS 0.1711 0.2512 - - 0.0541 0.0876 - -
FMGS Zuo2024FMGS 0.0916 0.2028 0.1169 0.2684 0.0641 0.1374 0.0579 0.1443
GOI QU2024GOI 0.2295 0.4638 0.2115 0.4286 0.1193 0.2297 0.1001 0.2240
@@
Gradient-Weighted 3DGS joseph2024GWfeatureBP 0.4184 0.5174 0.3573 0.4638 0.1638 0.2502 0.0601 0.1100
LUDVIG marrie2024ludvig 0.1406 0.2885 0.1337 0.2969 0.0887 0.1833 0.0602 0.1701
OccamLGS cheng2024occam 0.3559 0.4587 0.2308 0.4148 0.1936 0.2458 0.1204 0.2503
SceneSplat li2025SceneSplat7k (Pseudo Label) 0.3521 0.5210 0.3500 0.5550 0.2520 0.3038 0.2280 0.3590
@@
SceneSplat li2025SceneSplat7k 0.2968 0.5711 0.3540 0.5780 0.1387 0.4198 0.1648 0.3573
Table 3: Zero-shot 3D Semantic Segmentation Experiments on ScanNet20 dai2017scannet (20 classes) and ScanNet200 dai2017scannet (200 classes). All methods are evaluated on a 10-scene mini-validation set, with the 312-scene full validation set run only for selected methods due to runtime limitations (Tab.˜1).
@@
LangSplat Qin2024LangSplat 0.0131 0.0332 - - 0.0219 0.1257 - -
OpenGaussian Wu2024OpenGaussians 0.0695 0.1186 - - 0.1387 0.2698 - -
Feature-3DGS zhou2024feature3DGS 0.0798 0.1765 - - 0.3043 0.4732 - -
FMGS Zuo2024FMGS 0.0984 0.1892 0.1004 0.2300 0.2105 0.3664 - -
GOI QU2024GOI 0.1593 0.2790 0.1631 0.3296 0.1631 0.3296 0.1652 0.3130
@@
Gradient-Weighted 3DGS joseph2024GWfeatureBP 0.0925 0.1391 0.0900 0.1328 0.3083 0.4586 0.2762 0.4187
LUDVIG marrie2024ludvig 0.0995 0.2099 0.1099 0.2547 0.2348 0.3888 0.1960 0.3738
OccamLGS cheng2024occam 0.1580 0.2861 0.1502 0.3312 0.1803 0.3263 0.1725 0.3151
SceneSplat li2025SceneSplat7k (Pseudo Label) 0.2340 0.3709 0.2243 0.4667 0.2392 0.4368 0.2748 0.4384
@@
SceneSplat li2025SceneSplat7k 0.2263 0.4916 0.2836 0.4992 0.2732 0.5379 0.3384 0.5745
Table 4: Zero-Shot 3D Semantic Segmentation Experiments On ScanNet++ yeshwanth2023scannet++ and Matterport3D chang2017matterport3d. All methods are evaluated on a 10-scene mini-validation set, with the full validation set evaluated only for selected methods due to runtime limitations (Tab.˜1).
   Method    ScanNet20 (10 scenes)    ScanNet200 (10 scenes)    ScanNet++ (10 scenes)
   mIoU    mAcc    mIoU    mAcc    mIoU    mAcc
   Per‑Scene Optimization Methods
   LangSplat Qin2024LangSplat    0.0193    0.0700    0.0133    0.0272    0.0111    0.0272
   OpenGaussian Wu2024OpenGaussians    0.1958    0.3294    0.0740    0.1273    0.1290    0.2133
   Feature-3DGS zhou2024feature3DGS    0.5054    0.6474    0.1867    0.2798    0.1605    0.2368
   FMGS Zuo2024FMGS    0.1413    0.2629    0.0848    0.1825    0.1134    0.2244
   GOI QU2024GOI    0.2866    0.4536    0.1134    0.1806    0.1635    0.2767
   Per‑Scene Optimization‑Free Methods
   Gradient-Weighted 3DGS joseph2024GWfeatureBP    0.4533    0.5785    0.1639    0.2367    0.1210    0.1760
   LUDVIG marrie2024ludvig    0.1673    0.3058    0.0958    0.1828    0.1239    0.2588
   OccamLGS cheng2024occam    0.2937    0.4645    0.1300    0.2146    0.1670    0.2866
   Generalizable Method
   SceneSplat li2025SceneSplat7k    0.4458    0.6727    0.2599    0.4049    0.3165    0.5094
Table 5: Zero-Shot 2D Semantic Segmentation Experiments on ScanNet20 dai2017scannet (20 classes), ScanNet200 dai2017scannet (200 classes) and ScanNet++ yeshwanth2023scannet++. All methods are evaluated on a 10-scene validation set.
   Method    HoliCity (10 scenes)    Holicity (328 scenes)
   f–mIoU    f–mAcc    f–mIoU    f–mAcc
   Per‑Scene Optimization Methods
   LangSplat Qin2024LangSplat    0.0918    0.1950    -    -
   OpenGaussian Wu2024OpenGaussians    0.1853    0.2567    -    -
   Feature‑3DGS zhou2024feature3DGS    0.2479    0.5646    -    -
   FMGS Zuo2024FMGS    0.1774    0.2513    -    -
   GOI QU2024GOI    0.1126    0.3773    0.1587    0.3772
   Per‑Scene Optimization‑Free Methods
   Gradient-Weighted 3DGS joseph2024GWfeatureBP    0.2484    0.4808    0.2762    0.4187
   LUDVIG marrie2024ludvig    0.1593    0.2296    0.1843    0.2775
   OccamLGS cheng2024occam    0.2255    0.4754    0.2234    0.5056
   SceneSplat li2025SceneSplat7k (Pseudo Label)    0.2464    0.4780    0.2522    0.5159
   Generalizable Method
   SceneSplat li2025SceneSplat7k    0.3081    0.5647    0.2880    0.6045
Table 6: Zero-Shot 3D Outdoor Segmentation On HoliCity zhou2021holicity Dataset. All methods are evaluated on mini-validation sets,with the full validation set evaluated only for selected methods due to runtime limitations (Tab.˜1).
   Method    ScanNet (10 scenes)    ScanNet++ (10 scenes)
   Accuracy (BBox)    Accuracy (Seg)    Accuracy (BBox)    Accuracy (Seg)
   Per-Scene Optimization Methods
   LangSplat Qin2024LangSplat    0.0663    0.0415    0.0655    0.0337
   OpenGaussian Wu2024OpenGaussians    0.4565    0.3483    0.5138    0.3394
   Feature-3DGS zhou2024feature3DGS    0.5861    0.4759    0.5046    0.2783
   FMGS Zuo2024FMGS    0.6082    0.4639    0.5494    0.3303
   GOI QU2024GOI    0.5269    0.4414    0.5107    0.3568
   Per-Scene Optimization-Free Methods
   Gradient-Weighted 3DGS joseph2024GWfeatureBP    0.5486    0.4669    0.4322    0.2661
   LUDVIG marrie2024ludvig    0.6750    0.5481    0.4801    0.2875
   OccamLGS cheng2024occam    0.2074    0.1599    0.1896    0.1182
   Generalizable Method
   SceneSplat li2025SceneSplat7k    0.7701    0.6928    0.7339    0.5464
Table 7: 3D Object Localization Experiments on ScanNet dai2017scannet and ScanNet++ yeshwanth2023scannet++. Accuracy is reported with bounding box–based and segmentation–based evaluation.
@@
Table 8: Impact of Training-Data Scaling on Indoor Benchmarks and Cross-Domain Generalization to HoliCity zhou2021holicity. Results show that more training data consistently improves indoor performance. Furthermore, models trained on indoor data only surprisingly transfer to outdoor scenes.